OpenStudy (anonymous):

The sum of the squares of 3 consecutive positive integers is 116. What are the numbers? Which of the following equations is used in the process of solving this problem? 3n2 + 5 = 116 3n2 + 3n + 3 = 116 3n2 + 6n + 5 = 116

OpenStudy (anonymous):

If you let the number as n, then other two consecutive numbers will be: n+1, n + 2 So: \[n^2 + (n+1)^2 + (n+2)^2 = 116\] Just solve this..

OpenStudy (anonymous):

To open up the squares you may use this: \[(x+y)^2 = x^2 + y^2 + 2xy\]

OpenStudy (anonymous):

Idk how to do this

OpenStudy (anonymous):

Can you solve this: \[(n + 1)^2\]

OpenStudy (anonymous):

Just show me how to slove it

OpenStudy (anonymous):

You just need to sustitute it in the formula (a+b)^2=a^2+2ab+b^2..can u do it now??.. Just take n as a and 1as b and substitute the equation @waterineyes has provided... Okay??